### v1

# The initial value problem in cosmology: An alternative derivation for beginners

Abstract
In this article we address the puzzle related to the existence of a proper cosmological solution, where the number of equations exceeds the number of unknown functions. Although only two of the various dynamical equations are independent, the solution of those two equations must have to satisfy all the dynamical equations simultaneously at all instants. We show that this last requirement demands a constraint on the initial condition. When the initial constraint is satisfied, then and only then, one can have proper cosmological dynamics. The whole exercise turns out to be an alternative derivation of the initial value problem in cosmology.
1 Introduction
A first course in cosmology is rarely preceded by an advanced course in general relativity (GR). In an advanced course in general relativity, one first learns about time like, spacelike or null hypersurfaces. The geometry of hypersurfaces naturally demands the concepts of the first fundamental form and the second fundamental form of the hypersurfaces. Based on the last two conceptions one then proceeds toward the initial value problem in cosmology where the initial conditions are given on a spacelike 3-dimensional hypersurface. To understand the formulation of the initial value problem of the Einstein equation, in the cosmological backdrop, one requires to have a fair amount of understanding of differential geometry and advanced tensor analysis. There are some very good books that can be used by the advanced readers to get hold of these subtle issues [1, 2]. Most of the introductory books on cosmology, as in Ref. [3, 4, 5] do not address these issues as these issues are not that important for understanding cosmological models and the various phenomenological aspects of cosmology. The other books which primarily focus on gravitation, as in Ref. [6, 7], and discuss cosmology as purely an application of GR does not address these issues in any detail. There are a good number of papers related to the dynamics of the Friedmann-Lemaitre-Robertson-Walker (FLRW) model in cosmology, as Refs. [8, 9, 10, 11, 12], where the authors discuss various aspects of cosmology. These sets of papers are written for beginners and are both informative as well as quantitative. None of these literature does discuss the initial value problem of the Einstein equations. Most of the authors comment on the interdependence of the various dynamical equations and then start to analyze the various cosmological models resulting from various kinds of matter present in the universe. This is a perfect way to start and the authors cannot be blamed for not including the initial value problem because the initial value problem, for the case of Einstein equation, is more mathematics-oriented. A good discussion on this issue can be obtained in Ref. [13], but this reference is not suitable for readers who have not taken a course in differential geometry and topology. As a result of the complexity arising from the very nature of the Einstein equation, the initial value problem remains untouched by beginners who start to understand cosmology from university courses. In this article, we will try to address this initial value problem, in the cosmological case. Although the problem is mathematical in nature we will try to present the substance in such a way so that even students who have some knowledge of the Friedmann equations will be able to grasp the main points.
In FLRW cosmology it is known that there are multiple dynamical equations that are obtained by various kinds of combinations of the two Einstein equations. The combination of the two equations may not always be purely algebraic, one may need to differentiate one of the equations and add the result suitably with the other. As a result of these combinations, one can obtain four different equations (or more) [^1] The number of equations may depend upon the scheme in which the two Einstein equations are manipulated and various authors work with various kinds of those schemes. In this article we will primarily focus on four equations. The arguments presented in this article can be applied for any other schemes. out of which only two are independent. It turns out that there are two unknown functions: one is the scale-factor appearing in the FLRW metric and the other is the energy density of matter. Here we are assuming that the fluid representing matter has a barotropic equation of state. In such a case we have more number of equations than unknowns. Although the four equations are not independent, still they are different equations. For a proper solution of the Einstein equation, all of the four differential equations must have to agree simultaneously for all relevant time instants when the actual dynamics is produced from only two differential equations. We will show that a unique dynamics is only obtained when one chooses the initial condition judiciously. The initial conditions, in the case of the cosmological solutions, are constrained by one of the Einstein equations. This constraint plays a vital role in cosmological dynamics. The results we obtain are similar to the one described in Ref. [1] and as a result, our work can be interpreted as an alternative derivation of the cosmological initial value problem. Although the initial value problem in GR is primarily a mathematical problem, it has immense importance to the cosmologists who routinely work with Friedmann equations. If one is not fully aware of the initial value problem one can easily produce wrong results unknowingly.
The initial value problem for the Einstein equation has been worked out previously and we do not claim to present any original new result. We claim that our work is a new alternative derivation of the cosmological initial value problem which can be appreciated by anyone who has just started to know cosmology and has taken an introductory course in GR. The standard derivation of the cosmological initial value problem is inherently more geometric in nature where the initial conditions are specified on a spacelike 3-dimensional Cauchy surface. The effect of the (3+1)-dimensional Einstein equation produces constraints on the initial value on the Cauchy surface. The standard derivation is mathematically rigorous and complete. It has only one disadvantage, it cannot be accessed by students who have not taken some courses on differential geometry or advanced GR. In contrast, our alternative derivation is less geometric in nature. In deriving our results we use the continuity of the cosmological solution and time symmetry of the Einstein equation. Throughout this article, we have emphasized more on intuition. We think the new students who have just started to learn Friedmann equations will be able to follow the whole article.
The material in the article is presented in the following way. In the next section, we present the cosmological solution of the Einstein equation. In this section, we have not emphasized the derivation of the equations as we assume the reader to know these results. This section primarily sets the convention followed in the rest of the article. In section 3, we have posed the question related to the dynamics emanating from the Einstein equation. Section 4 tries to answer the previous question for ever-expanding or ever-contracting spacetimes. In this section, it is seen that the results may depend upon the existence of a turning point, where the time derivative of the scale-factor of the FLRW spacetime changes sign. In section 5 we address the problems related to the turning points. The last section concludes the article with a brief discussion of the results.
2 A single barotropic fluid in FLRW spacetime and the Einstein equation
Let us start from the beginning. We start with the Einstein equation in general relativity (GR):
$G_{\mu\nu}\equiv R_{\mu\nu}-\frac{1}{2}Rg_{\mu\nu}=8\pi GT_{\mu\nu}+\Lambda g_{\mu\nu}\,,$ (1)
where $g_{\mu\nu}$ is the metric tensor, $R_{\mu\nu}$ is the Ricci tensor, $R=g^{\mu\nu}R_{\mu\nu}$ is the Ricci scalar, $T_{\mu\nu}$ is the energy-momentum tensor for matter, $\Lambda$ is the cosmological constant and $G$ is the universal gravitational constant. In our units, the speed of light $c=1$. The affine connection coefficients are specified as
$\Gamma^{\mu}_{\alpha\beta}=\frac{1}{2}g^{\mu\sigma}(\partial_{\alpha}g_{\sigma\beta}+\partial_{\beta}g_{\sigma\alpha}-\partial_{\sigma}g_{\alpha\beta})\,.$ (2)
The Riemann tensor is given by
$R^{\mu}_{\,\,\,\,\nu\alpha\beta}=\partial_{\alpha}\Gamma^{\mu}_{\nu\beta}-\partial_{\beta}\Gamma^{\mu}_{\nu\alpha}+\Gamma^{\mu}_{\sigma\alpha}\Gamma^{\sigma}_{\nu\beta}-\Gamma^{\mu}_{\sigma\beta}\Gamma^{\sigma}_{\nu\alpha}\,.$ (3)
Here we are using the Einstein summation convention and any repeated index is summed up. In the above equation we have two summations on $\sigma$. Henceforth we will always use this summation convention. We get the Ricci tensor from the Riemann tensor by contraction of suitable indices as
$R_{\nu\beta}\equiv R^{\mu}_{\,\,\,\,\nu\mu\beta}=\partial_{\mu}\Gamma^{\mu}_{\nu\beta}-\partial_{\beta}\Gamma^{\mu}_{\nu\mu}+\Gamma^{\mu}_{\sigma\mu}\Gamma^{\sigma}_{\nu\beta}-\Gamma^{\mu}_{\sigma\beta}\Gamma^{\sigma}_{\nu\mu}\,.$ (4)
Standard textbooks on GR specifies that due to the Bianchi identity, followed by the Riemann tensor, one will always have
$\nabla_{\mu}G^{\mu\nu}=0\,,$ (5)
where the covariant derivative is defined as
$\nabla_{\mu}G^{\mu\nu}=\partial_{\mu}G^{\mu\nu}+\Gamma^{\mu}_{\mu\alpha}G^{\nu\alpha}+\Gamma^{\nu}_{\mu\alpha}G^{\mu\alpha}\,.$ (6)
The above result is important as it sets an important condition on the energy-momentum tensor for matter.
In GR the affine connection coefficients are symmetric in the lower indices. The Ricci tensor and the metric tensor are also symmetric second rank tensors. This implies $G_{\mu\nu}$ is a symmetric tensor and consequently $T_{\mu\nu}$ should also be a symmetric tensor. From Eq. (5) we immediately get
$\nabla_{\mu}T^{\mu\nu}=0\,.$ (7)
The original Einstein equation remains consistent as because in GR the metric tensor is covariantly constant: $\nabla_{\mu}g^{\mu\nu}=0$. It is seen that the conservation of the energy-momentum tensor is a condition arising from the structure of the Einstein tensor.
The spacetime metric is chosen to be the isotropic and homogeneous FLRW spacetime:
$ds^{2}=-dt^{2}+a^{2}(t)\left[\frac{dr^{2}}{1-kr^{2}}+r^{2}(d\theta^{2}+\sin^{2}\theta d\phi^{2})\right]\,,$ (8)
where $a(t)$ is the scale-factor for the expanding or contracting spacetime and $k$ is a constant which can take values ±$\pm 1$ or 00. The spatial coordinate $r$, used in the above line element is called a comoving coordinate. One must be careful in interpreting the above line element. Previously we have said that we are using units where speed of light is one. As a consequence length and time have the same dimension. We can interpret $ds^{2}$ having a dimension of squared length. In this unit what will be the dimensions of $r$, $a(t)$ and $k$? When $k=0$, $r$ does have the dimension of length and $a(t)$ can be dimensionless. When $k=\pm 1$ and dimensionless, $r$ cannot carry length dimension. In such cases $a(t)$ carries the dimension of length. As $k$ comes out from the geometry of the FLRW spacetime, we will like to keep it dimensionless.
In this work we will assume that the FLRW spacetime is seeded by a single component barotropic fluid whose energy momentum tensor is given by
$T_{\mu\nu}=(\rho+P)u_{\mu}u_{\nu}+Pg_{\mu\nu}\,,$ (9)
where $\rho$ is the energy density of the fluid, $P$ is its pressure and the fluid ‘elements move’ with a time like 4-velocity $u_{\mu}=dx^{\mu}/d\tau$ where
$u_{\mu}u^{\mu}=-1\,.$
In FLRW coordinates the proper time, $\tau$, is the same as the coordinate time $t$ itself. In the FLRW metric, in the comoving coordinate, we will have $u^{\mu}=(1,{\bf 0})$ where ${\bf 0}$ specifies that the fluid elements do not move spatially. For a barotropic fluid one must also have to specify the equation of state (EoS):
$P=\omega\rho\,,$ (10)
where $\omega$, which is a dimensionless quantity in our system of units, is often called the EoS. The EoS is a constant. For a perfect fluid the only nonzero components of the energy-momentum tensor are
$T_{00}=\rho\,,\,\,\,\,\,\,T_{ij}=Pg_{ij}\,.$ (11)
Written in this form one knows how the energy-momentum tensor components are related with the energy density and pressure of the fluid.
In FLRW spacetime, the nonzero components of the Einstein tensor are $G_{00}$ and $G_{ij}$ and these tensor components can be equated to the only nonzero components of the energy-momentum tensor $T_{00}$ and $T_{ij}$. The relevant equations are $G_{00}=8\pi GT_{00}+\Lambda g_{00}$ and $G_{ij}=8\pi GT_{ij}+\Lambda g_{ij}$. These two equations yield
$\frac{\dot{a}^{2}}{a^{2}}+\frac{k}{a^{2}}$ $=$ $\frac{8\pi G}{3}\rho+\frac{\Lambda}{3}\,,$ (12)
$2\frac{\ddot{a}}{a}+\frac{\dot{a}^{2}}{a^{2}}+\frac{k}{a^{2}}$ $=$ $-8\pi GP+\Lambda\,,$ (13)
where the dot specifies a derivative with respect to time, $t$. The first of the above two equations is called the second Friedmann equation by many authors. Subtracting Eq. (12) from Eq. (13) one gets
$\frac{\ddot{a}}{a}=-\frac{4\pi G}{3}(\rho+3P)+\frac{\Lambda}{3}\,.$ (14)
Many authors call the above equation to be the first Friedmann equation. The last equation is also called the acceleration equation. In this work we will call Eq. (12) as the Friedmann equation, the second equation, Eq. (13), will be called as the pressure equation and the last equation will be named as the acceleration equation.
Except the above three equations, which govern the dynamics of the system, there is another one. This equation comes from the energy-momentum conservation condition given in Eq. (7). In FLRW spacetime this condition becomes
$\dot{\rho}+3\frac{\dot{a}}{a}(\rho+P)=0\,.$ (15)
As because this equation is an outcome of the fact that the left hand side of the Einstein equation satisfies the condition arising from Bianchi identity and is true because of the particular form of the Einstein tensor, this equation can be derived from the first Friedmann equation and the second Friedmann equation. We will call Eq. (15) as the continuity equation in this work. As like the continuity equations in other branches of physics, this equation also says how the energy density of the system continuously adjusts itself so that the net energy-momentum of the system is conserved.
Now we have four dynamical equations for the development of the FLRW spacetime. We emphasize here that, although we know that we have two independent differential equations coming out from the Einstein equation (the Friedmann equation and the the pressure equation), they cannot be simultaneously solved. The reason being, none of them contain derivatives of $\rho$ or $P$. In such a case one is bound to work with the continuity equation as it is the only one which has a time derivative of $\rho$. We have three functions to determine: $a(t)$, $\rho(t)$ and $P(t)$. Using the EoS we know $P(t)$ can be expressed in terms of the energy density and hence we have only two unknown functions. How do we figure out the development of this dynamical system? Of course all four of these equations are not independent of each other, only two of these four equations must be enough to determine the dynamical evolution of the system. How to choose which two of them are really relevant? We will answer this question in a subsequent section. In the next section we specify the problem in its full generality.
3 Setting up the dynamical system
To specify the problem in its full generality we first give a simple definition of a dynamical system. A dynamical system consists of a set of differential equations and a specific initial configuration. Given these two things the dynamical system is expected to predict the time development of the system. In our case we have a dynamical system at hand. Our dynamical system consists of four ordinary, nonlinear, coupled differential equations of different order and a set of initial conditions. The problem is that, we have four differential equations and we require only two out of them and consequently two of the equations will be left out from our system. Suppose the dynamical system evolves from time $t_{i}$ to $t_{f}$ via the joint solution of the two chosen differential equations. What will happen to the other two differential equations which were left out? If the system is not an overdetermined system, consistency requires that the solutions $a(t)$ and $\rho(t)$, during evolution from $t_{i}$ to $t_{f}$, must also satisfy the two equations which are left out during the same time interval. As all the four equations available to us are valid differential equations expressing some physical content, but may not be independent of each other, all of them must agree for any arbitrary time, $t$, in the interval $t_{i}\leq t\leq t_{f}$. One has to be a bit careful to show the agreement as all the differential equations are not of the same degree, some are first order differential equations and some are second order differential equations. As the various differential equations are not of the same degree one has to find out a way to specify equivalent initial conditions for all the differential equations. We will discuss the point of equivalent initial conditions later.
We can now be more specific about the problem of time evolution in the FLRW system. Is it always possible that the four differential equations agree during $t_{i}\leq t\leq t_{f}$ while the actual dynamics is produced by only two differential equations satisfying some specific initial conditions, keeping $a(t)>0$ and $\rho(t)>0$ throughout the time interval $[t_{i}\,,\,t_{f}]$? The answer turns out to be positive for a certain set of initial conditions. For initial conditions which are outside the preferred set of initial conditions, all the four equations will never match and the system will behave as an overdetermined system. In the last case we have to conclude that the Einstein equation does not yield any specific solution. We present various cases to illustrate the various possibilities in the present problem.
To present our arguments we will first work with cosmologies involving purely expanding or contracting phases. Later we will discuss the cases where an expanding (contracting) phase can convert to a contracting (expanding) phase.
4 The various cases in an ever expanding or ever contracting universe
In this section we will discuss about purely expanding or contracting phases in cosmology. Throughout the relevant interval, $t_{i}\leq t\leq t_{f}$, we assume that $\dot{a}\neq 0$. We will relax the last assumption in the next section and show that for intervals where we have $\dot{a}(t_{0})=0$, at some point $t_{0}$ in the interval, we can still obtain specific solutions of the Einstein equation. To present our case systematically we present a chronological description of the dynamics produced by various pairs of the relevant equations. We will first try to see whether we can reproduce the differential equations left out from our chosen pair of equations. Although the four differential equations are produced essentially from two independent dynamical equations, it will be seen that one may not get back all the equations from any pair of relevant equations. We will then try to see whether the solution produced from the chosen pair of equations simultaneously satisfy the other equations.
Before we start we must point out that the time interval $[t_{i}\,,\,t_{f}]$ has to be chosen carefully. We have demanded the interval to be such that the scale-factor always increases or decreases with time inside the interval. We have tacitly assumed that $a(t)>0$ inside the interval. The last condition may not be true for any time interval. For a contracting spacetime, the dynamics can produce $a(t_{s})=0$ for some $t_{i}\leq t_{s}\leq t_{f}$. In such cases the system has reached a singularity. In this case one has to redefine the interval in such a way that the upper limit of the relevant interval, $t_{f}$, always remains smaller than $t_{s}$. In other words, our interval must be chosen in such a way that there are no singularities inside this interval.
4.1 Choosing the Friedmann equation and the continuity equation
When we choose the Friedmann equation and the continuity equation, we see that both of them are first order differential equations. They require two independent initial conditions: $a(t_{i})$ and $\rho(t_{i})$. We have not chosen the pressure equation and the acceleration equation in our scheme. We will show that if $\dot{a}\neq 0$, then in general the pressure equation is obtained from our chosen equations. When we solve the Friedmann equation and the continuity equation we do not explicitly require the expression of $\ddot{a}$. If we try to figure out the value of $\ddot{a}$ from our chosen equations, we exactly get the pressure equation.
To obtain the value of $\ddot{a}$ we first take a time derivative of both the sides of the Friedmann equation and get
$\frac{\dot{a}}{a}\left(\frac{2\ddot{a}}{a}-\frac{2\dot{a}^{2}}{a^{2}}-\frac{2k}{a^{2}}\right)=\frac{8\pi G}{3}\dot{\rho}=-8\pi G\frac{\dot{a}}{a}(\rho+P)\,,$ (16)
where we have utilized the continuity equation to write the last step. Assuming $\dot{a}\neq 0$ in the interval $t_{i}\leq t\leq t_{f}$ we get
$\frac{2\ddot{a}}{a}-\frac{2\dot{a}^{2}}{a^{2}}-\frac{2k}{a^{2}}=-8\pi G(\rho+P)\,.$ (17)
Writing the above equation as
$\frac{2\ddot{a}}{a}-\frac{2\dot{a}^{2}}{a^{2}}-\frac{2k}{a^{2}}+8\pi G\rho=-8\pi GP\,,$ (18)
and then utilizing the Friedmann equation, to replace $8\pi G\rho$, we get the pressure equation. This shows that the values of $a(t)$ and $\rho(t)$ generated from the Friedmann equation also satisfies the pressure equation for any $t$ inside the interval $[t_{i}\,,\,t_{f}]$ for cases where $\dot{a}\neq 0$ inside the time interval. Now one can again combine the pressure equation with the Friedmann equation to obtain the acceleration equation. As a consequence of this the generated values of $a(t)$ and $\rho(t)$ will also satisfy the acceleration equation. Hence all the relevant dynamical equations are satisfied simultaneously in this case.
4.2 Choosing the pressure equation and the continuity equation
In this scheme, we choose the pressure equation and the continuity equation as our basic dynamical equations. Here one of the equations is a second order differential equation, while the other one is a first order differential equation. Using the continuity equation we write the pressure equation as
$2\frac{\ddot{a}}{a}+\frac{\dot{a}^{2}}{a^{2}}+\frac{k}{a^{2}}=8\pi G\left(\frac{1}{3}\dot{\rho}\frac{a}{\dot{a}}+\rho\right)+\Lambda\,,$ (19)
where we have explicit appearance of $\dot{\rho}/\dot{a}$. The above equation always remain well defined as because $\dot{a}$ never vanishes in the time interval $[t_{i}\,,\,t_{f}]$. Multiplying all the terms of the above equation by $\dot{a}/a$ we get
$2\frac{\ddot{a}\dot{a}}{a^{2}}+\frac{\dot{a}^{3}}{a^{3}}+\frac{k\dot{a}}{a^{3}}=\frac{8\pi G}{3}\dot{\rho}+8\pi G\rho\frac{\dot{a}}{a}+\Lambda\frac{\dot{a}}{a}\,.$ (20)
The above equation can also be written as
$\frac{d}{dt}\left(\frac{\dot{a}^{2}}{a^{2}}\right)+\frac{3\dot{a}^{3}}{a^{3}}+\frac{k\dot{a}}{a^{3}}=\frac{8\pi G}{3}\dot{\rho}+8\pi G\rho\frac{\dot{a}}{a}+\Lambda\frac{\dot{a}}{a}\,.$
Subtracting
$3\left(\frac{\dot{a}^{3}}{a^{3}}+\frac{k\dot{a}}{a^{3}}\right)$
from both sides of the last equation, we get
$\frac{d}{dt}\left(\frac{\dot{a}^{2}}{a^{2}}+\frac{k}{a^{2}}\right)=\frac{8\pi G}{3}\frac{d\rho}{dt}+3\frac{\dot{a}}{a}\left[\frac{8\pi G}{3}\rho+\frac{\Lambda}{3}-\left(\frac{\dot{a}^{2}}{a^{2}}+\frac{k}{a^{2}}\right)\right]\,.$ (21)
At this stage, this equation does not convey anything interesting. It can become interesting if we define
$F(t)\equiv\frac{8\pi G}{3}\rho+\frac{\Lambda}{3}-\left(\frac{\dot{a}^{2}}{a^{2}}+\frac{k}{a^{2}}\right)\,,$ (22)
and write Eq. (21) as
$\dot{F}+3\frac{\dot{a}}{a}F=0\,.$ (23)
This equation shows that the Friedmann equation may not hold when we obtain the dynamics of the system from the pressure equation and the continuity equation. If and only if $F(t)=0$ for all instants in the interval $[t_{i}\,,\,t_{f}]$, the Friedmann equation will be valid throughout the interval. In general for arbitrary initial conditions, $F(t)$ is never zero at any point inside the interval. For any initial condition $a(t_{i})$, $\dot{a}(t_{i})$ and $\rho(t_{i})$ we can always solve the continuity equation to get a positive $\rho(t)$ for the cases where $\omega=0$ or $\omega=1/3$. Using the expression of $\rho(t)$ in the pressure equation we can get $a(t)$. Although we have produced the dynamics from an arbitrary, but otherwise valid, initial condition, this dynamical evolution will not satisfy the Friedmann equation. The resultant dynamical evolution cannot be a solution to the Einstein equation.
4.3 Choosing the acceleration equation and the continuity equation
In this case we start with the acceleration equation and the continuity equation. This choice of equations is interesting as because none of the equations in our scheme have any dependence on $k$. Looking at the equations one cannot guess anything about the spatial curvature of the FLRW spacetime. Nonetheless, these equations are capable of producing dynamical evolution of the system when we impose reasonable initial conditions: $a(t_{i})$, $\dot{a}(t_{i})$ and $\rho(t_{i})$. As because our dynamical system lacks any information about $k$, it may seem that it is impossible to obtain the other differential equations from these two equations as the other equations require specification of $k$. We will opine on this interesting question in the next subsection.
To proceed, we first rewrite the acceleration equation as
$\frac{\ddot{a}}{a}=-\frac{4\pi G}{3}(\rho+P)-\frac{8\pi G}{3}P+\frac{\Lambda}{3}\,.$ (24)
Using the continuity equation, we can write the last equation as
$\frac{\ddot{a}}{a}=\frac{4\pi G}{9}\frac{a}{\dot{a}}\dot{\rho}-\frac{8\pi G}{3}P+\frac{\Lambda}{3}\,.$ (25)
This equation is also well defined throughout the interval $[t_{i}\,,\,t_{f}]$ as $\dot{a}$ never vanishes inside the interval. Multiplying both sides of the above equation by $2\dot{a}/a$ we get
$\frac{2\dot{a}\ddot{a}}{a^{2}}=\frac{8\pi G}{9}\dot{\rho}-\frac{16\pi G}{3}\frac{\dot{a}}{a}P+\frac{2\dot{a}\Lambda}{3a}\,.$ (26)
In the next step we will subtract
$2\frac{\dot{a}}{a}\left(\frac{\dot{a}^{2}}{a^{2}}+\frac{\tilde{k}}{a^{2}}\right)\,,$
from both sides of the above equation. At this level, $\tilde{k}$, is simply an arbitrary constant, whose value is not yet specified. Subtracting the aforementioned quantity from both sides of the last equation and manipulating some terms we get
$\frac{d}{dt}\left(\frac{\dot{a}^{2}}{a^{2}}+\frac{\tilde{k}}{a^{2}}\right)=\frac{8\pi G}{3}\dot{\rho}-\frac{16\pi G}{9}\dot{\rho}-\frac{16\pi G}{3}\frac{\dot{a}}{a}P+\frac{2\dot{a}\Lambda}{3a}-2\frac{\dot{a}}{a}\left(\frac{\dot{a}^{2}}{a^{2}}+\frac{\tilde{k}}{a^{2}}\right)\,.$ (27)
The above equation can be written as
$\frac{d}{dt}\left(\frac{\dot{a}^{2}}{a^{2}}+\frac{\tilde{k}}{a^{2}}-\frac{8\pi G}{3}\rho-\frac{\Lambda}{3}\right)=-\frac{16\pi G}{9}\left(\dot{\rho}+\frac{3\dot{a}}{a}P\right)+\frac{2\dot{a}\Lambda}{3a}-2\frac{\dot{a}}{a}\left(\frac{\dot{a}^{2}}{a^{2}}+\frac{\tilde{k}}{a^{2}}\right)\,.$ (28)
Defining $\tilde{F}(t)$ as
$\tilde{F}(t)\equiv\frac{8\pi G}{3}\rho+\frac{\Lambda}{3}-\left(\frac{\dot{a}^{2}}{a^{2}}+\frac{\tilde{k}}{a^{2}}\right)\,,$ (29)
and using the continuity equation, we get
$\dot{\tilde{F}}$ $=$ $-\frac{16\pi G}{3}\frac{\dot{a}}{a}\rho-\frac{2\dot{a}\Lambda}{3a}+2\frac{\dot{a}}{a}\left(\frac{\dot{a}^{2}}{a^{2}}+\frac{\tilde{k}}{a^{2}}\right)$ (30)
$=$ $-2\frac{\dot{a}}{a}\tilde{F}\,,$
which implies
$\dot{\tilde{F}}+2\frac{\dot{a}}{a}\tilde{F}=0\,.$ (31)
This equation is similar to Eq. (23) but not exactly the same. The last equation is difficult to interpret as we still cannot fix the value of $\tilde{k}$ from our chosen set of equations. In such a case we cannot even guess whether the Friedmann equation is valid. Even we do not know from our set of equations whether the pressure equation is valid. For some arbitrary initial conditions we can generate dynamical evolution from our set of equations but we will not be able to know whether our solution satisfies the Friedmann equation and the pressure equation. The situation is more serious than the case before. If we start with any arbitrary initial conditions we do not even know whether the Einstein equation is valid.
4.4 The link between all the choices is in the initial condition
From our previous discussion we have seen that all the pairs of chosen differential equations may not reproduce the other differential equations. In such cases one can have a resultant dynamics which is not a unique solution of the Einstein equation. All the previous discussion was based on arbitrary choice of initial conditions. We did not mention any scheme for comparing initial conditions. As because our system admits both first order and second order differential equations, it is difficult to give an ansatz for equivalence of initial condition. Equivalence of initial condition means an initial condition which produces a proper solution to the Einstein equation, no matter whichever pair of equations one works with. When we use such an initial condition, we demand all the chosen pairs of differential equations will produce the same time development of the system and all the differential equations will be simultaneously satisfied at all instants of time inside the chosen interval $[t_{i}\,,\,t_{f}]$. The pressure equation and continuity equation as well as the acceleration equation and the continuity equation requires three numbers in the initial condition: $a(t_{i})$, $\dot{a}(t_{i})$ and $\rho(t_{i})$. There is no unique way to choose this triplet of numbers. On the other hand the Friedmann equation and the continuity equation only requires two initial conditions: $a(t_{i})$ and $\rho(t_{i})$. Apparently it seems it is impossible to specify the initial condition which can be applied to all the three pairs of equations. If we treat all the differential equations as purely dynamical equations then it is really impossible to find out an equivalent initial condition.
The above problem can be tackled if we reinterpret the Friedmann equation as a constraint equation at the initial point $t_{i}$. The Friedmann equation is the only equation which which can be interpreted as an algebraic equation connecting the values of $\dot{a}(t_{i})$, $a(t_{i})$ and $\rho(t_{i})$ for some values of $k$ and $\Lambda$. From our previous discussion we have seen that when we produce the dynamics from the pressure equation and the continuity equation from any arbitrary initial condition the Friedmann equation may not hold. But once we choose the initial conditions from the Friedmann equation itself, using it as a constraint equation, we see that the Friedmann equation remains valid throughout the time interval $[t_{i}\,,\,t_{f}]$. This happens because of the nature of Eq. (23). If the initial condition satisfies the Friedmann constraint then $F(t_{i})=0$ and consequently $\dot{F}(t_{i})=0$. As a consequence $F(t)$ is always zero inside the interval $[t_{i}\,,\,t_{f}]$. This is a particularly interesting phenomena. The initial constraint condition which produces the initial values of the variables, becomes a dynamical entity and, remains dynamically true throughout the relevant time interval. The constraint propagates through time. As we chose the pressure equation and the continuity equation to generate the dynamics, we see that if the initial condition is chosen from the Friedmann constraint then the Friedmann equation remains valid throughout the time interval. The generated dynamics is naturally a solution of the Friedmann equation. Once the initial condition is chosen from the Friedmann constraint, the resulting dynamical solution satisfies all the relevant dynamical equations, simultaneously, in $[t_{i}\,,\,t_{f}]$.
Next we focus our attention to the dynamics generated by the acceleration equation and the continuity equation. In this case we do not know the value of $\tilde{k}$. If the initial condition is chosen from the Friedmann equation, with a specific value of $k$, then what happens? We see from Eq. (29), this initial condition can only be chosen when $\tilde{k}=k$. In this case, the Friedmann constraint gives the information about the spatial curvature of the initial 3-dimensional hypersurface. Not only the initial condition fixes the value of $k$, now $\tilde{F}(t)=F(t)$, where $F(t)$ is defined as in Eq. (22). In this case also we see, from Eq. (29), that if $\tilde{F}(t_{i})=0$ then $\tilde{F}(t)=0$, at all instants in the interval $[t_{i}\,,\,t_{f}]$. It is interesting to note that although the chosen pair of dynamical equations did not have any information about $k$, appearing in the Friedmann equation, the initial condition passes on the information and this information is not lost as the initial constraint propagates through time. The result shows that if we choose the initial condition through the Friedmann constraint then the Friedmann equation always remains true. All the four relevant dynamical differential equations will have the same solution if we choose the initial condition via the Friedmann constraint.
From the previous discussion it is seen that the Friedmann equation acts as a constraint equation, based on which the initial condition of cosmological dynamics is obtained. This constraint equation remains valid throughout the interval $[t_{i}\,,\,t_{f}]$ and consequently it can also be interpreted as a dynamical equation. The dual roles do not contradict because one can choose any time instant to be the initial time $t_{i}$. As a consequence the Friedmann equation becomes a dynamic constraint.
5 The cases where $\dot{a}(t_{0})=0$ and $t_{0}$ is inside the interval $[t_{i}\,,\,t_{f}]$
The instant when $\dot{a}$ vanishes, is an instant when an expanding spacetime starts to contract or it can also be an instant, when an initially contracting spacetime starts to expand. The first case is called gravitational collapse and the second case is called a cosmological bounce, respectively. In these cases one has to interpret the results carefully. The first point to note is that both in Eq. (19) and in Eq. (25) we have terms where $\dot{\rho}/\dot{a}$ appear explicitly. Apparently this terms may diverge when $\dot{a}$ vanishes. The fact is, none of these terms diverge at any moment inside the relevant time interval. At the instant when $\dot{a}$ vanishes, $\dot{\rho}$ also vanishes and their ratio remains finite. From the continuity equation we get
$\frac{\dot{\rho}}{\dot{a}}=-\frac{3(\rho+P)}{a}\,.$
Even when both $\dot{\rho}$ and $\dot{a}$ tends to zero, inside the interval $[t_{i}\,,\,t_{f}]$, their ratio remains finite.
5.1 When $t_{0}=t_{i}$
In this case we have $\dot{a}(t_{i})=0$ at the initial point. The Friedmann equation and the continuity equation can generate the dynamics from this point. The discussion in subsection 4.1 showed us that the Friedmann equation and the continuity equation naturally implies the pressure equation as long as $\dot{a}(t)\neq 0$ inside $[t_{i}\,,\,t_{f}]$. Although the pressure equation connot be connected to the Friedmann equation and the continuity equation at $t=t_{i}$, one can still have a specific solution to the full Einstein equation. That a solution exists in the present case can be shown in the following way. We know that the Friedmann equation and the continuity equation can generate a continuous $a(t)$ and $\rho(t)$ in the time interval $[t_{i}\,,\,t_{f}]$. Suppose instead of $t_{i}$, we start at $t_{i}+\epsilon$, where $\epsilon$ is an arbitrarily small positive constant with dimension of time. We assume that $\dot{a}(t_{i})=0$ and consequently $\dot{a}$ never vanishes inside the interval $[t_{i}+\epsilon\,,\,t_{f}]$. The initial condition at $t=t_{i}+\epsilon$ are now generated from the known solution coming out from the Friedmann equation and continuity equation, which gave the dynamical solution in the full time interval $[t_{i}\,,\,t_{f}]$. All the generated solutions for different (but arbitrarily small) $\epsilon$ smoothly matches with the original solution defined in the full time interval for regions $t_{i}+\epsilon\leq t\leq t_{f}$. From our discussion in subsection 4.2 we must conclude that inside the interval $[t_{i}+\epsilon\,,\,t_{f}]$ we must have a proper solution to the Einstein equation as the pressure equation and the continuity equation will also have the same solution produced from the Friedmann equation and continuity equation if the initial condition is obtained from the Friedmann constraint. Now one can make $\epsilon$ arbitrarily small and for each small value of $\epsilon$ we always obtain a unique cosmological solution. Continuity of our generated solution implies that in the limit $\epsilon\to 0$ we will also have a unique solution of the Einstein equation. The solutions from the two different pairs coincide at $t_{i}=t_{0}$ as the pressure equation and the continuity equation starts with initial conditions following the Friedmann constraint.
For the other cases related to Eq. (23) and Eq. (31), we see that if $\dot{a}(t_{i})=0$, it seems that both these equations can be satisfied for any $F(t_{i})$ as long as $\dot{F}(t_{i})=0$. But this arbitrariness disappears once we use the Friedmann constraint initially and fix $F(t_{i})=0$. As a result of this the Friedmann equation will be valid for all points inside the relevant time interval and we will always get a definite solution to the Einstein equation.
5.2 When $t_{i}<t_{0}<t_{f}$
Next we discuss the case where $\dot{a}(t_{0})=0$ where $t_{0}$ is an intermediate point in the relevant time interval. In this case we can obtain a solution to the Einstein equation if our initial condition satisfies the Friedmann constraint. We can obtain the collapsing solution in the interval $[t_{i}\,,\,t_{f}]$ by using the Friedmann equation and the continuity equation. But is this solution also obtained by the other pairs of equations? Before we proceed we observe that all the relevant dynamical equations, namely the Friedmann equation, pressure equation, acceleration equation and the continuity equation are symmetric around $t=0$, by which we mean
$a(t)=a(-t)\,,\,\,\,\,\rho(t)=\rho(-t)\,,\,\,\,\,\dot{a}(t)=-\dot{a}(-t)\,,\,\,\,\,\ddot{a}(t)=\ddot{a}(-t)\,,\,\,(t\geq 0)\,,$ (32)
[…]
6 Conclusion
In this section we summarize what we have discussed till now. It is seen that in the FLRW spacetime, seeded by matter and a cosmological constant, we obtain two independent dynamical equations from the Einstein equation. These equations are nonlinear differential equations for the scale-factor $a(t)$. One of these equations is a first order differential equation while the other one is a second order differential equation. These two dynamical equations in general is incapable of producing the dynamics of the system as they do not contain and time derivative of the energy density or pressure. In such a case one has to include the continuity equation in the list of workable equations. The continuity equation is not an independent differential equation, it can be derived from the previous two equations. One can also combine the Einstein equations to obtain another second order differential equation called the acceleration equation. We have now four differential equations and two unknown functions: $a(t)$ and $\rho(t)$. The pressure is related to the energy density via the EoS of a barotropic fluid. As the number of equations exceeds the number of unknowns do we have an overdetermined system of differential equations? This is the primary question which is addressed in this article.
To answer the question properly one has to first observe that the system of equations we are dealing with are not algebraic equations: they are ordinary nonlinear differential equations. Differential equations by themselves do not give enough information about the existence of a unique solution, one also has to include proper initial conditions. It is observed that for the Einstein equation, in the cosmological backdrop, the initial condition plays a vital role. As only two differential equations, out of four, are enough for determining the dynamics of the system one has to be careful about the fate of the two differential equations left out. If the solution is an unique solution then it must simultaneously satisfy all the four dynamical equations. It is seen that different pairs of dynamical equations do not encode the same amount of information. The pair consisting of the Friedmann equation and the continuity equation only require two initial conditions as they are both first order differential equations. This pair is capable of producing the dynamics coming out of the Einstein equation. This shows that the system requires only two independent initial conditions. On the other hand the other pairs of dynamical equations each contain one second order differential equation. These systems require more information as they require three initial conditions. The natural question which arises is how can the different dynamics which are produced from the various pairs of differential equations be the same if they carry different amounts of information?
In this article we have tried to answer the question raised in the last paragraph. When one tries to answer the question one discovers that actually one of the Einstein equations, the Friedmann equation namely, plays a dual role. It itself can produce the dynamical evolution of the system in conjunction with the continuity equation. On the other hand it also acts as a constraint equation from which we obtain the initial condition. One will get an unique cosmological dynamics from any pair of differential equations when the initial condition is obtained from the Friedmann equation acting as a constraint equation. More interestingly as the pairs which contain a second order differential equation contains more information than required, the Friedmann equation acts as a dynamical constraint. Even if our choice of differential equations do not involve the Fredmann equation but the initial condition is obtained from the Friedmann constraint, the Friedmann equation constrains the dynamics in the relevant time interval of our interest. As a consequence we see that even for second order differential equations, their degrees of freedom remains restricted and the initial constraint propagates through time. The system does not become an overdetermined system if the initial condition is chosen using the Friedmann constraint. If the initial condition does not respect the Friedmann constraint one can still obtain a dynamics of the system from the pairs of equations containing second order differential equations, but in this case the solutions will not be unique.
All the results presented in this article concerns the dynamical evolution of the FLRW spacetime where the matter part is made up of a barotropic fluid. It is interesting to note that most of the results presented in this article also holds when the matter part is made up of a scalar field. There will be some minor changes in the results as now one has to redefine the initial conditions. For the sake of completeness we present briefly the results for the scalar field in Appendix B.
The primary question posed in the article and its answer provided above constitute the cosmological initial value problem in GR. We claim that we have presented an alternative derivation of the initial value problem in cosmology. Our method is general. Although we have worked with only one fluid one can easily include more than one fluid components in the above analysis.
Acknowledgement: DD would like to acknowledge the support of the Atlantic Association for Research in the Mathematical Sciences (AARMS) for funding the work.
Appendix
Appendix A Shifting of origin in the cosmological setup
In section 5, we commented that the shifting of the origin in the case of the FLRW cosmological solutions is a nontrivial process. In this appendix we show why it is so. We know that Eq. (32) is valid when the origin of time is at $t=0$. Suppose in the cosmological backdrop we know the evolution of the system in the time interval $[t_{i}\,,\,t_{f}]$ To make the arguments straightforward we can assume that the system is expanding in the relevant time interval and $t_{i}>0$. In this case the initial condition is applied at $t=t_{i}$. Can someone shift the origin to some time $\tilde{t}$, where $t_{i}<\tilde{t}<t_{f}$, for the aforementioned expansion process? If one shifts the origin to $\tilde{t}$, the relevant time interval changes to $[t_{i}-\tilde{t}\,,\,t_{f}-\tilde{t}]$ and the initial condition is applied at $t=t_{i}-\tilde{t}$. The answer to the last question is no, as because in the interval where we have expansion, $\dot{a}(\tilde{t})\neq 0$. We see from Eq. (32) that $\dot{a}(0)=0$ and consequently if $\dot{a}$ does not vanish at some point of time we will never be able to assume that point of time to be the origin and claim that the new dynamics in the interval $[t_{i}-\tilde{t}\,,\,t_{f}-\tilde{t}]$ is simply obtained from the old dynamics in the interval $[t_{i}\,,\,t_{f}]$ by a shift in the origin of time.
If one applies the initial condition to the Einstein equation, in FLRW cosmology, at some time $t=0$ and is only interested to know about the future development of the system or the past development of the system then one may choose $\dot{a}(0)\neq 0$. The resultant dynamics will be unique if one uses the Friedmann constraint. On the other hand if there is a past and future of $t=0$ and $\dot{a}(0)\neq 0$, and the initial condition is applied at $t=0$, then the system will become inconsistent and one will never be able to find out the proper cosmological dynamics from the Einstein equation.
Things are different for the case where $\dot{a}(t_{0})=0$ for some $t_{0}$ in the interval $[t_{i}\,,\,t_{f}]$. In this case one can actually shift the origin of time to $t_{0}$ without any difficulty. In bouncing cosmologies or gravitational collapse one can always shift the origin of time to $t_{0}$ when the turn over happens. In this case one will get the same bounce or collapse in the shifted interval $[t_{i}-t_{0}\,,\,t_{f}-t_{0}]$. The reason why this happens is because one can now properly give the initial condition at $t=0$. The time shifted dynamics respects the time symmetry of the Einstein equation even at $t=0$ because $\dot{a}(t_{0})=0$ The resulting bounce or gravitational collapse will always be symmetric with respect to the moment $t=t_{0}$.
In cosmology, one often studies gravitational collapse of an isolated spherical overdensity of energy, decoupled from the cosmological expansion of spacetime, in the late time universe. In the latter case one starts (the initial expanding phase) at some $t=0$ where $\dot{a}(0)\neq 0$. In this case the system remains well defined as one never asks the question what happened to the spherical overdensity before $t=0$. It is tacitly assumed that the overdensity started to grow at $t=0$ and ultimately collapsed to produce a virialized halo structure later. To know more about this collapsing processes one may read about the simplest gravitational collapse, named Top Hat collapse, from Ref. [3].
Appendix B Cosmology with a canoncal scalar field
Till now we have assumed that the matter part is represented by a barotropic fluid. Most of the results stated above remain true even when we have a canonical scalar field instead of a barotropic fluid. There are some changes from the previous results and to point them out we discuss the case of scalar field cosmology briefly in this section. By canonical scalar field we mean a scalar field whose kinetic term is given by the conventional kinetic term for scalar fields.
As the FLRW spacetime is assumed to be isotropic and homogeneous, the scalar field, $\phi(t)$, can only be a function of the cosmic time. The scalar field dynamics requires the specification of the scalar field potential, $V(\phi)$. For a scalar field, $\phi(t)$, the barotropic equation of state does not work as its energy density and pressure are given by
$\rho_{\phi}=\frac{\dot{\phi}^{2}}{2}+V(\phi)\,,\,\,\,\,\,P_{\phi}=\frac{\dot{\phi}^{2}}{2}-V(\phi)\,.$ (33)
All the four equations: the Friedmann equation, pressure equation, acceleration equation, and continuity equation are also applicable in the present case. The continuity equation now becomes
$\dot{\rho}_{\phi}+3\frac{\dot{a}}{a}(\rho_{\phi}+P_{\phi})=0\,,$ (34)
which gives
$\dot{\phi}\left(\ddot{\phi}+3\frac{\dot{a}}{a}\dot{\phi}+\frac{dV}{d\phi}\right)=0\,.$
From the above equation, it may seem that at the moment when $\dot{\phi}=0$ any value of $\phi$ will satisfy the energy conservation equation. In reality the the value of $\phi$ at that moment is not arbitrary. This is because although the velocity for the scalar field may be zero at a point of time, it still has an acceleration term, $\ddot{\phi}$, at that moment. To properly define the acceleration term in the time interval $[t_{i}\,,\,t_{f}]$ the continuity equation is interpreted as a field equation:
$\ddot{\phi}+3\frac{\dot{a}}{a}\dot{\phi}+\frac{dV}{d\phi}=0\,,$ (35)
irrespective of the value of $\dot{\phi}$. Although the continuity equation and the above field equation are equivalent, the field equation contains more information about the scalar field. In practice one has to solve the field equation to obtain the values of $\dot{\phi}$ and $\phi$ which defines $\rho_{\phi}$ and $P_{\phi}$. In this case the initial condition consists of $a(t_{i})$, $\dot{a}(t_{i})$, $\phi(t_{i})$ and $\dot{\phi}(t_{i})$. For a proper solution of the Einstein equation the initial condition must satisfy the Friedmann constraint as well as the scalar field equation at $t=t_{i}$.
All the results discussed in this article hold true for the scalar fields as well, one has to use $\rho_{\phi}$ and $P_{\phi}$ in the place of fluid energy density and pressure. The equation of continuity is as given in Eq. (34). Formally the similarity ends there. There is one big difference between barotropic fluids and scalar fields. In the case of barotropic fluid, the energy density was purely evolving via the Einstein equation. In the case of the scalar field, the energy density and pressure also evolve via the Einstein equation but the dynamics of the scalar field are also constrained by the form of the potential $V(\phi)$.
The potential is a function of $\phi(t)$ and hence the potential is a function of time. This potential function may be symmetric in time or may not be symmetric. Depending upon the initial condition and symmetry of the potential, the Einstein equation may be symmetric or asymmetric about the point $t=0$, even when $\dot{a}(0)=0$. For some potentials and for some specific initial conditions, one may have $\rho_{\phi}(-t)\neq\rho_{\phi}(t)$. In these cases Eq. (32) does not hold even when $\dot{a}(0)=0$. In such cases, one may find asymmetric cosmological bounce. For the symmetric bounce, the initial value problem is tackled in the way we tackled symmetric gravitational collapse in this article. For the asymmetric bounce, one can still shift the origin of time at the bounce time, $t=t_{0}$, and apply the initial condition, using the Friedmann constraint, at that point of time. The dynamics of the system, both in the future and the past of this point, can be generated from any pair of dynamical equations. This dynamics will satisfy all the differential equations. If one wants one can now shift back the origin of time to $t=t_{i}$.