### v2

# Redundancy of the cosmological evolution equations and its relationship with the initial conditions

Abstract
It is known that in Friedmann-Lemaitre-Robertson-Walker cosmology one has more number of dynamical equations, compared to the number of unknown variables. This fact makes some equations redundant. The situation becomes complicated because all the relevant differential equations in cosmology are not of the same order. In this article we study the fate of the redundant equations. We show that this redundancy is inevitable in general relativity and a modern modification of gravity, called $f(R)$ gravity. It is shown that this redundancy predicts a special role to one of the Friedmann equations, which constrains the initial values of the problem. It is shown that the initial constraint condition itself evolves with time. Our method of analyzing the dynamical structure of the theories relies on an operational approach and can be generalized further.
1 Introduction
The basics of modern cosmology are taught at the masters level in various institutions. The standard paradigm of introductory cosmology is well addressed in Refs. [1, 2, 3] whereas Refs. [4, 5, 6] focusses more on general relativity (GR). These books give a general understanding based on which someone can build up his/her conceptions in cosmology. Readers who will like to know more about the various aspects of modern cosmology based on the Friedmann-Lemaitre-Robertson-Walker (FLRW) spacetime can consult the excellent pedagogical articles in Refs. [7, 8, 9, 10, 11]. While going through the basics of cosmology one first learns to tackle the dynamical equations in cosmology, commonly called the Friedmann equations. These Friedmann equations are nonlinear ordinary differential equations which originate from the Einstein equations in GR which are second order nonlinear partial differential equations in three spatial and one time dimensions. While dealing with these equations one first observes that even in the simplest situation, where the universe only contains one barotropic fluid, there is a problem of redundancy. There are simply more number of cosmological evolution equations than required. The situation is more challenging as all the available differential equations are not of the same order. Naturally the question arises why such a problem appears at all and how are we going to obtain proper cosmological development from these equations in a consistent manner. In this article we will primarily be addressing this redundancy problem in cosmology. We will first show that this redundancy problem is intimately related to the initial value problem in cosmology. It turns out that any arbitrary initial condition will not produce proper cosmological dynamics, only initial conditions satisfying a particular constraint will produce proper dynamics. Then we show that the redundancy problem in cosmology has its own logic in GR. We will also briefly state how our analysis in GR can be generalized in a more modern variation of a theory of gravity, called $f(R)$ gravity.
The article is intended for readers who have taken a basic course of introductory cosmology and knows how to solve the Friedmann equations in standard scenarios: As in matter domination or in radiation domination [^1] We assume that the intended reader to be aware of at least some of the references cited in the last paragraph.. What has been presented in this article has deeper connections to the theory of second order nonlinear partial differential equations, involving time development, in three dimensional space.
The interesting point which becomes apparent from the present article is related to the fact that we can learn many interesting properties related to cosmological development in GR, as well as in $f(R)$ gravity, from an operational approach where our aim is solely to solve the dynamical equations unambiguously. Such an approach ultimately leads to a thorough understanding of the logical structure of the theories.
2 A single barotropic fluid in FLRW spacetime and the Einstein equation
The spacetime metric, in cosmology, is the isotropic and homogeneous FLRW spacetime:
$ds^{2}=-dt^{2}+a^{2}(t)\left[\frac{dr^{2}}{1-kr^{2}}+r^{2}(d\theta^{2}+\sin^{2}\theta d\phi^{2})\right]\,$, (1)
where $a(t)$ is the scale-factor for the expanding or contracting spacetime and $k$ is a constant which can take values $\pm 1$ or 0. If $k=0$, the curvature of the spatial hypersurfaces in FLRW spacetime is zero, indicating flat spatial hypersurfaces. On the other hand k=+1,1k=+1,\,-1, specifies that the spatial hypersurfaces are positively curved (as on the surface of a sphere) or negatively curved (as in a saddle), respectively. The spatial coordinate $r$, used in the above line element is called a comoving coordinate. One must be careful in interpreting the above line element. We are using units where speed of light is one. As a consequence length and time have the same dimension. We can interpret $ds^{2}$ having a dimension of squared length. In this unit what will be the dimensions of $r$, $a(t)$ and $k$? When $k=0$, $r$ does have the dimension of length and $a(t)$ can be dimensionless. When $k=\pm 1$ and dimensionless, $r$ cannot carry length dimension. In such cases $a(t)$ carries the dimension of length. As $k$ comes out from the geometry of the FLRW spacetime, we will like to keep it dimensionless.
In this work we will assume that the FLRW spacetime is seeded by a single component barotropic fluid whose equation of state (EoS):
$P=\omega\rho\,$, (2)
where $\omega$, which is a dimensionless quantity in our system of units, is often called the EoS parameter. Here $\rho$ is the energy-density of the fluid and $P$ specifies its isotropic pressure. The EoS parameter is a constant. In presence of a barotropic fluid, the Einstein equations of FLRW cosmology are:
$\frac{\dot{a}^{2}}{a^{2}}+\frac{k}{a^{2}}$ $=$ $\frac{8\pi G}{3}\rho+\frac{\Lambda}{3}\,$, (3)
$2\frac{\ddot{a}}{a}+\frac{\dot{a}^{2}}{a^{2}}+\frac{k}{a^{2}}$ $=$ $-8\pi GP+\Lambda\,$, (4)
where the dot specifies a derivative with respect to time, $t$. The first of the above two equations is called the second Friedmann equation by many authors. Subtracting Eq. (3) from Eq. (4) one gets
$\frac{\ddot{a}}{a}=-\frac{4\pi G}{3}(\rho+3P)+\frac{\Lambda}{3}\,.$ (5)
Many authors call the above equation to be the first Friedmann equation. The last equation is also called the acceleration equation. In this work we will call Eq. (3) as the Friedmann equation, the second equation, Eq. (4), will be called as the pressure equation and the last equation will be named as the acceleration equation.
Except the above three equations, which govern the dynamics of the system, there is another one. This equation comes from the energy-momentum conservation condition. The energy-momentum conservation condition in GR is an outcome of the Bianchi identity. In FLRW spacetime this condition becomes
$\dot{\rho}+3\frac{\dot{a}}{a}(\rho+P)=0\,.$ (6)
We will call Eq. (6) as the continuity equation in this work. As like the continuity equations in other branches of physics, this equation also says how the energy density of the system continuously adjusts itself so that the net energy-momentum of the system is conserved.
Now we have four dynamical equations for the development of the FLRW spacetime. We emphasize here that, although we know that we have two independent differential equations coming out from the Einstein equation (the Friedmann equation and the the pressure equation), they cannot be simultaneously solved. The reason being, none of them contain derivatives of $\rho$ or $P$. In such a case one is bound to work with the continuity equation as it is the only one which has a time derivative of $\rho$. We have three functions to determine: $a(t)$, $\rho(t)$ and $P(t)$. Using the EoS we know $P(t)$ can be expressed in terms of the energy density and hence we have only two unknown functions. How do we figure out the development of this dynamical system?
We can now be more specific about the problem of time evolution in the FLRW system. Is it always possible that the four differential equations agree during $t_{i}\leq t\leq t_{f}$ while the actual dynamics is produced by only two differential equations satisfying some specific initial conditions, keeping $a(t)>0$ and $\rho(t)>0$ throughout the time interval $t_{i}\,,\,t_{f}]$? The answer turns out to be positive for a certain set of initial conditions. For initial conditions which are outside the preferred set, all the four equations will never match and the system will behave as an overdetermined system. In the last case we have to conclude that the Einstein equation does not yield any specific solution. We present various cases to illustrate the various possibilities in the present problem.
To present our arguments we will first work with cosmologies involving purely expanding or contracting phases. Later we will discuss the cases where an expanding (contracting) phase can convert to a contracting (expanding) phase.
3 The various cases in an ever expanding or ever contracting universe
In this section we will discuss about purely expanding or contracting phases in cosmology. Throughout the relevant interval, $t_{i}\leq t\leq t_{f}$, we assume that $\dot{a}\neq 0$. We will relax the last assumption in the next section and show that for intervals where we have $\dot{a}(t_{0})=0$, at some point $t_{0}$ in the interval, we can still obtain specific solutions of the Einstein equation. We will first try to see whether we can reproduce the differential equations left out from our chosen pair of equations. Although the four differential equations are produced essentially from two independent dynamical equations, it will be seen that one may not get back all the equations from any pair of relevant equations. We will then try to see whether the solution produced from the chosen pair of equations simultaneously satisfy the other equations.
Before we start we must point out that the time interval $t_{i}\,,\,t_{f}]$ has to be chosen carefully. We have demanded the interval to be such that the scale-factor always increases or decreases with time inside the interval. We have tacitly assumed that $a(t)>0$ inside the interval. The last condition may not be true for any time interval. For a contracting spacetime, the dynamics can produce $a(t_{s})=0$ for some $t_{i}\leq t_{s}\leq t_{f}$. In such cases the system has reached a singularity. In this case one has to redefine the interval in such a way that the upper limit of the relevant interval, $t_{f}$, always remains smaller than $t_{s}$. In other words, our interval must be chosen in such a way that there are no singularities inside this interval.
3.1 Choosing the Friedmann equation and the continuity equation
When we choose the Friedmann equation and the continuity equation, we see that both of them are first order differential equations. They require two independent initial conditions: $a(t_{i})$ and $\rho(t_{i})$. We have not chosen the pressure equation and the acceleration equation in our scheme. To obtain the value of $\ddot{a}$ we first take a time derivative of both the sides of the Friedmann equation and get
$\frac{\dot{a}}{a}\left(\frac{2\ddot{a}}{a}-\frac{2\dot{a}^{2}}{a^{2}}-\frac{2k}{a^{2}}\right)=\frac{8\pi G}{3}\dot{\rho}=-8\pi G\frac{\dot{a}}{a}(\rho+P)\,$, (7)
where we have utilized the continuity equation to write the last step. Assuming $\dot{a}\neq 0$ in the interval $t_{i}\leq t\leq t_{f}$ we get
$\frac{2\ddot{a}}{a}-\frac{2\dot{a}^{2}}{a^{2}}-\frac{2k}{a^{2}}=-8\pi G(\rho+P)\,.$ (8)
Writing the above equation as
$\frac{2\ddot{a}}{a}-\frac{2\dot{a}^{2}}{a^{2}}-\frac{2k}{a^{2}}+8\pi G\rho=-8\pi GP\,$, (9)
and then utilizing the Friedmann equation, to replace $8\pi G\rho$, we get the pressure equation. This shows that the values of $a(t)$ and $\rho(t)$ generated from the Friedmann equation also satisfies the pressure equation for any $t$ inside the interval $t_{i}\,,\,t_{f}]$ for cases where $\dot{a}\neq 0$ inside the time interval. Now, one can again combine the pressure equation with the Friedmann equation to obtain the acceleration equation. As a consequence of this the generated values of $a(t)$ and $\rho(t)$ will also satisfy the acceleration equation. Hence all the relevant dynamical equations are satisfied simultaneously in this case.
3.2 Choosing the pressure equation and the continuity equation
In this scheme, we choose the pressure equation and the continuity equation as our basic dynamical equations. Here one of the equations is a second order differential equation, while the other one is a first order differential equation. Using the continuity equation we write the pressure equation as
$2\frac{\ddot{a}}{a}+\frac{\dot{a}^{2}}{a^{2}}+\frac{k}{a^{2}}=8\pi G\left(\frac{1}{3}\dot{\rho}\frac{a}{\dot{a}}+\rho\right)+\Lambda\,$, (10)
where we have explicit appearance of $\dot{\rho}/\dot{a}$. The above equation always remain well defined as because $\dot{a}$ never vanishes in the time interval $t_{i}\,,\,t_{f}]$. Multiplying all the terms of the above equation by $\dot{a}/a$ we get
$2\frac{\ddot{a}\dot{a}}{a^{2}}+\frac{\dot{a}^{3}}{a^{3}}+\frac{k\dot{a}}{a^{3}}=\frac{8\pi G}{3}\dot{\rho}+8\pi G\rho\frac{\dot{a}}{a}+\Lambda\frac{\dot{a}}{a}\,.$ (11)
The above equation can also be written as
$\frac{d}{dt}\left(\frac{\dot{a}^{2}}{a^{2}}\right)+\frac{3\dot{a}^{3}}{a^{3}}+\frac{k\dot{a}}{a^{3}}=\frac{8\pi G}{3}\dot{\rho}+8\pi G\rho\frac{\dot{a}}{a}+\Lambda\frac{\dot{a}}{a}\,.$
Subtracting
$3\left(\frac{\dot{a}^{3}}{a^{3}}+\frac{k\dot{a}}{a^{3}}\right)$
from both sides of the last equation, we get
$\frac{d}{dt}\left(\frac{\dot{a}^{2}}{a^{2}}+\frac{k}{a^{2}}\right)=\frac{8\pi G}{3}\frac{d\rho}{dt}+3\frac{\dot{a}}{a}\left[\frac{8\pi G}{3}\rho+\frac{\Lambda}{3}-\left(\frac{\dot{a}^{2}}{a^{2}}+\frac{k}{a^{2}}\right)\right]\,.$ (12)
At this stage, this equation does not convey anything interesting. It can become interesting if we define
$F(t)\equiv\frac{8\pi G}{3}\rho+\frac{\Lambda}{3}-\left(\frac{\dot{a}^{2}}{a^{2}}+\frac{k}{a^{2}}\right)\,$, (13)
and write Eq. (12) as
$\dot{F}+3\frac{\dot{a}}{a}F=0\,.$ (14)
This equation shows that the Friedmann equation may not hold when we obtain the dynamics of the system from the pressure equation and the continuity equation. If and only if $F(t)=0$ for all instants in the interval $t_{i}\,,\,t_{f}]$, the Friedmann equation will be valid throughout the interval. In general for arbitrary initial conditions, $F(t)$ is never zero at any point inside the interval. Consequently for arbitrary initial conditions this dynamical evolution will not satisfy the Friedmann equation. The resultant dynamical evolution cannot be a valid solution of the cosmological evolution equations.
3.3 Choosing the acceleration equation and the continuity equation
In this case we start with the acceleration equation and the continuity equation. This choice of equations is interesting as because none of the equations in our scheme have any dependence on $k$. Looking at the equations one cannot guess anything about the spatial curvature of the FLRW spacetime. Nonetheless, these equations are capable of producing dynamical evolution of the system when we impose reasonable initial conditions: $a(t_{i})$, $\dot{a}(t_{i})$ and $\rho(t_{i})$. As because our dynamical system lacks any information about $k$, it may seem that it is impossible to obtain the other differential equations from these two equations as the other equations require specification of $k$. We will opine on this interesting question in the next subsection.
To proceed, we first rewrite the acceleration equation as
$\frac{\ddot{a}}{a}=-\frac{4\pi G}{3}(\rho+P)-\frac{8\pi G}{3}P+\frac{\Lambda}{3}\,.$ (15)
Using the continuity equation, we can write the last equation as
$\frac{\ddot{a}}{a}=\frac{4\pi G}{9}\frac{a}{\dot{a}}\dot{\rho}-\frac{8\pi G}{3}P+\frac{\Lambda}{3}\,.$ (16)
This equation is also well defined throughout the interval $t_{i}\,,\,t_{f}]$ as $\dot{a}$ never vanishes inside the interval. Multiplying both sides of the above equation by $2\dot{a}/a$ we get
$\frac{2\dot{a}\ddot{a}}{a^{2}}=\frac{8\pi G}{9}\dot{\rho}-\frac{16\pi G}{3}\frac{\dot{a}}{a}P+\frac{2\dot{a}\Lambda}{3a}\,.$ (17)
In the next step we will subtract
$2\frac{\dot{a}}{a}\left(\frac{\dot{a}^{2}}{a^{2}}+\frac{\tilde{k}}{a^{2}}\right)\,$,
from both sides of the above equation. At this level, $\tilde{k}$, is simply an arbitrary constant, whose value is not yet specified. Subtracting the aforementioned quantity from both sides of the last equation and manipulating some terms we get
$\frac{d}{dt}\left(\frac{\dot{a}^{2}}{a^{2}}+\frac{\tilde{k}}{a^{2}}\right)=\frac{8\pi G}{3}\dot{\rho}-\frac{16\pi G}{9}\dot{\rho}-\frac{16\pi G}{3}\frac{\dot{a}}{a}P+\frac{2\dot{a}\Lambda}{3a}-2\frac{\dot{a}}{a}\left(\frac{\dot{a}^{2}}{a^{2}}+\frac{\tilde{k}}{a^{2}}\right)\,.$ (18)
The above equation can be written as
$\frac{d}{dt}\left(\frac{\dot{a}^{2}}{a^{2}}+\frac{\tilde{k}}{a^{2}}-\frac{8\pi G}{3}\rho-\frac{\Lambda}{3}\right)=-\frac{16\pi G}{9}\left(\dot{\rho}+\frac{3\dot{a}}{a}P\right)+\frac{2\dot{a}\Lambda}{3a}-2\frac{\dot{a}}{a}\left(\frac{\dot{a}^{2}}{a^{2}}+\frac{\tilde{k}}{a^{2}}\right)\,.$ (19)
Defining $\tilde{F}(t)$ as
$\tilde{F}(t)\equiv\frac{8\pi G}{3}\rho+\frac{\Lambda}{3}-\left(\frac{\dot{a}^{2}}{a^{2}}+\frac{\tilde{k}}{a^{2}}\right)\,$, (20)
and using the continuity equation, we get
$\dot{\tilde{F}}$ $=$ $-\frac{16\pi G}{3}\frac{\dot{a}}{a}\rho-\frac{2\dot{a}\Lambda}{3a}+2\frac{\dot{a}}{a}\left(\frac{\dot{a}^{2}}{a^{2}}+\frac{\tilde{k}}{a^{2}}\right)$ (21)
$=$ $-2\frac{\dot{a}}{a}\tilde{F}\,$,
which implies
$\dot{\tilde{F}}+2\frac{\dot{a}}{a}\tilde{F}=0\,.$ (22)
This equation is similar to Eq. (14) but not exactly the same. The last equation is difficult to interpret as we still cannot fix the value of $\tilde{k}$ from our chosen set of equations. In such a case we cannot even guess whether the Friedmann equation is valid.
3.4 The link between all the choices is in the initial condition
From our previous discussion we have seen that all the pairs of chosen differential equations may not reproduce the other differential equations. All the previous discussion was based on arbitrary choice of initial conditions. We did not mention any scheme for comparing initial conditions. The pressure equation and continuity equation as well as the acceleration equation and the continuity equation requires three numbers in the initial condition: $a(t_{i})$, $\dot{a}(t_{i})$ and $\rho(t_{i})$. There is no unique way to choose this triplet of numbers. On the other hand the Friedmann equation and the continuity equation only requires two initial conditions: $a(t_{i})$ and $\rho(t_{i})$. Apparently it seems it is impossible to specify the initial condition which can be applied to all the three pairs of equations. If we treat all the differential equations as purely dynamical equations then it is really impossible to find out an equivalent initial condition.
The above problem can be tackled if we reinterpret the Friedmann equation as a constraint equation at the initial point $t_{i}$. The Friedmann equation is the only equation which can be interpreted as an algebraic equation connecting the values of $\dot{a}(t_{i})$, $a(t_{i})$ and $\rho(t_{i})$ for some values of $k$ and $\Lambda$. Once we choose the initial conditions from the Friedmann equation itself, using it as a constraint equation, we see that the Friedmann equation remains valid throughout the time interval $t_{i}\,,\,t_{f}]$. This happens because of the nature of Eq. (14). If the initial condition satisfies the Friedmann constraint then $F(t_{i})=0$ and consequently $\dot{F}(t_{i})=0$. As we chose the pressure equation and the continuity equation to generate the dynamics, we see that if the initial condition is chosen from the Friedmann constraint then the Friedmann equation remains valid throughout the time interval. The generated dynamics is naturally a solution of the Friedmann equation.
Next we focus our attention to the dynamics generated by the acceleration equation and the continuity equation. In this case we do not know the value of $\tilde{k}$. If the initial condition is chosen from the Friedmann equation, with a specific value of $k$, then what happens? We see from Eq. (20), this initial condition can only be chosen when $\tilde{k}=k$. In this case, the Friedmann constraint gives the information about the spatial curvature of the initial 3-dimensional hypersurface. Not only the initial condition fixes the value of $k$, now $\tilde{F}(t)=F(t)$, where $F(t)$ is defined as in Eq. (13). In this case also we see, from Eq. (20), that if $\tilde{F}(t_{i})=0$ then $\tilde{F}(t)=0$, at all instants in the interval $t_{i}\,,\,t_{f}]$. All the four relevant dynamical differential equations will have the same solution if we choose the initial condition via the Friedmann constraint.
From the previous discussion it is seen that the Friedmann equation acts as a constraint equation, based on which the initial condition of cosmological dynamics is obtained. This constraint equation remains valid throughout the interval $t_{i}\,,\,t_{f}]$ and consequently it can also be interpreted as a dynamical equation. The dual roles do not contradict because one can choose any time instant to be the initial time $t_{i}$. As a consequence the Friedmann equation becomes a dynamic constraint.
4 The cases where $\dot{a}(t_{0})=0$ and $t_{0}$ is inside the interval $t_{i}\,,\,t_{f}]$
The instant when $\dot{a}$ vanishes, is an instant when an expanding spacetime starts to contract or it can also be an instant, when an initially contracting spacetime starts to expand. The first case is called gravitational collapse and the second case is called a cosmological bounce, respectively. In these cases one has to interpret the results carefully. The first point to note is that both in Eq. (10) and in Eq. (16) we have terms where $\dot{\rho}/\dot{a}$ appear explicitly. Apparently this terms may diverge when $\dot{a}$ vanishes. The fact is, none of these terms diverge at any moment inside the relevant time interval. At the instant when $\dot{a}$ vanishes, $\dot{\rho}$ also vanishes and their ratio remains finite. From the continuity equation we get
$\frac{\dot{\rho}}{\dot{a}}=-\frac{3(\rho+P)}{a}\,.$
Even when both $\dot{\rho}$ and $\dot{a}$ tends to zero, inside the interval $t_{i}\,,\,t_{f}]$, their ratio remains finite.
4.1 When $t_{0}=t_{i}$
In this case we have $\dot{a}(t_{i})=0$ at the initial point. That a solution exists in the present case can be shown in the following way. We know that the Friedmann equation and the continuity equation can generate a continuous $a(t)$ and $\rho(t)$ in the time interval $t_{i}\,,\,t_{f}]$. Suppose instead of $t_{i}$, we start at $t_{i}+\epsilon$, where $\epsilon$ is an arbitrarily small positive constant with dimension of time. We assume that $\dot{a}(t_{i})=0$ and consequently $\dot{a}$ never vanishes inside the interval $t_{i}+\epsilon\,,\,t_{f}]$. The initial conditions at $t=t_{i}+\epsilon$ are now generated from the known solution coming out from the Friedmann equation and continuity equation. All the generated solutions for different (but arbitrarily small) $\epsilon$ (possibly using other pairs of differential equations) smoothly matches with the solution defined in the full time interval for regions $t_{i}+\epsilon\leq t\leq t_{f}$. One can make $\epsilon$ arbitrarily small and continuity of our generated solution implies that in the limit $\epsilon\to 0$ we will also have a unique solution of the Einstein equation. The solutions from the different pairs coincide at $t_{i}=t_{0}$ as the pressure equation and the continuity equation starts with initial conditions following the Friedmann constraint.
For the other cases related to Eq. (14) and Eq. (22), we see that if $\dot{a}(t_{i})=0$, it seems that both these equations can be satisfied for any $F(t_{i})$ as long as $\dot{F}(t_{i})=0$. But this arbitrariness disappears once we use the Friedmann constraint initially and fix $F(t_{i})=0$. As a result of this the Friedmann equation will be valid for all points inside the relevant time interval (if we start from pairs of equations not containing the Friedmann equation) and we will always get a definite solution to the Einstein equation.
4.2 When $t_{i}<t_{0}<t_{f}$
Next we discuss the case where $\dot{a}(t_{0})=0$ where $t_{0}$ is an intermediate point in the relevant time interval. In this case we can obtain a solution to the Einstein equation if our initial condition satisfies the Friedmann constraint. We can follow a solution of the Einstein’s equation involving gravitational expansion and a subsequent gravitational contraction in the interval $t_{i}\,,\,t_{f}]$ by using the Friedmann equation and the continuity equation. The turn-around happens at $t=t_{0}$ when $\dot{a}(t_{0})=0$. Is this solution also obtained by the other pairs of equations? Before we proceed we observe that all the relevant dynamical equations, namely the Friedmann equation, pressure equation, acceleration equation and the continuity equation are symmetric around $t=0$, by which we mean
$a(t)=a(-t)\,,\,\,\,\,\rho(t)=\rho(-t)\,,\,\,\,\,\dot{a}(t)=-\dot{a}(-t)\,,\,\,\,\,\ddot{a}(t)=\ddot{a}(-t)\,,\,\,(t\geq 0)\,$, (23)
when $\dot{a}(t=0)=0$. We will use these properties and shift the origin of time to $t=t_{0}$ so that what used to be called $t_{0}$ becomes the origin of time $t=0$. The initial time interval now becomes $t_{i}-t_{0}\,,\,t_{f}-t_{0}]$. In shifted time we have $\dot{a}(0)=0$ which implies a symmetric gravitational collapse. The initial conditions can be well specified in the shifted time as $\rho(0)$ and $a(0)$. Now using any pair of differential equations we can generate the solution of the Einstein equation in the expanding phase, in the interval $t_{i}-t_{0},0]$. In this interval all the solutions will agree if all of them satisfy the Friedmann constraint initially. The symmetry of the Einstein equations now guarantee that these solutions must also agree on the contracting phase. If one now wants to go back to the old time interval, one has to just shift the time origin back to $t_{i}$.
5 On the logical consistency of the theory
From our previous discussion it is seen that one can model cosmological dynamics from any relevant pair of equations (presented in this article) and even if those equations do not involve the Friedmann equation specifically, one cannot forget about the Friedmann equation. All cosmological initial conditions must satisfy the Friedmann equation. Although we showed the relevance of this statement in a simple case, the last statement is valid in more general cases as will be shown later in this article.
The results of this article are related to the initial value problem and its solution in cosmology. Can one set up proper initial conditions in GR and solve the Einstein equations consistently in the future, where one is using the FLRW spacetime? This is the initial value problem in cosmology and it is known that the answer is in the affirmative. A formal discussion on the initial value problem in cosmology and its solution can be found in Refs. [12, 13]. In the rest of this section we discuss about the logical consistency of cosmological dynamics in GR.
Till now what has been discussed indicates that the complexity of the situation is primarily related to the availability of both first order and second order differential equations. Naturally the question arises: Is this a peculiarity of the cosmological equations in GR? In this section we will show that the dynamical structure of the cosmological theory, as predicted by GR, is logically consistent and makes perfect sense.
To understand the logical consistency in the present case, we will like to specify here that some form of a first order differential equation is inevitable in the present case if we have to make the theory meaningful. If we leave out quantum effects, we may safely assume that there cannot be matter creation from the gravitational field. In such a case we must have some condition which predicts that the evolution equation of energy density is a first order differential equation and its form must be $\dot{\rho}\propto\rho$. This is a very general requirement, as for such a relation if $\rho$ is zero initially then $\dot{\rho}$ is also zero initially and hence the matter energy density must remain zero for all times. A second order differential equation in $\rho(t)$ will never satisfy the general requirement, as in this case even if initially $\rho$ and $\dot{\rho}$ is zero, still $\rho(t)$ can attain finite non-zero values in the future.
As the Einstein equations (Friedmann and pressure equations) in cosmology do not contain a time derivative of the energy density of the fluid so those equations does not generate naturally a differential equation for $\rho(t)$, but we know from general assumptions, about the system, that such an equation is inevitable. In such a case there is only one way the first order differential equation for $\rho(t)$ can appear, that intended equation must be derivable from the available Einstein equations. In reality it happens like that as because the continuity equation is obtained from the Friedmann and pressure equations. To obtain the continuity equation one must have to differentiate the Friedmann equation and consequently the Friedmann equation must be a first order differential equation. The Friedmann equation cannot be a second order differential equation as in that case a differentiation of this equation will produce a third order derivative of the scale-factor in the resulting equation which cannot be cancelled by any terms in the other Einstein equation for pressure. Hence the equation giving the energy density of the fluid in terms of the functions of the scale-factor and its derivatives must have to be a first-order differential equation of $a(t)$. As because GR, as a well defined field theory, contains second order derivative terms of the scale-factor, the other equation for pressure must have to be a second order differential equation for $a(t)$.
At this point we must point out that with all the manipulations, as specified in the previous paragraphs, one in general always obtains the desired first order differential equation for $\rho(t)$. This is not a coincidence, rather the theory is formulated in such a way that the desired differential equation is produced due to a tensor identity, called the (second) Bianchi identity, constraining the Einstein tensor. The Bianchi identity naturally makes the theory consistent by introducing the energy conservation condition as predicted by the continuity equation.
The previous discussion made it clear that in GR, the Bianchi identity and the Einstein equation both are responsible for the dynamical structure of the cosmological theory and together they affect the initial condition of the problem. Interestingly, whatever we have discussed about cosmological dynamics in GR can be generalized to a modern modified theory of gravity. In the next section we generalize our results in $f(R)$ gravity.
6 On the redundancy of dynamical equations and initial conditions in $f(R)$ cosmology
Till now we have worked in the ambit of GR and Newtonian cosmology. In this section we will show that the techniques we have derived can be generalized in a modified theory of gravity, called the metric $f(R)$ theory. This modification of gravity is a modern development which became useful in theories of inflation and tackling the dark energy problem. The two reviews in Refs. [14, 15] will help the reader to grasp the essence of the theory and we will closely follow the conventions and notations used in those reviews.
In metric $f(R)$ gravity theory the gravitational action changes from
$S=\frac{1}{16\pi G}\int d^{4}x\sqrt{-g}\,R+S_{m}\to S=\frac{1}{16\pi G}\int d^{4}x\sqrt{-g}\,f(R)+S_{m}\,$,
where $R$ is the Ricci scalar and $S_{m}$ is the action for the non-gravitational degrees of freedom, generally called the matter action. Here $f(R)$ is a general function of $R$ and we do not choose any particular form of $f(R)$ here as our analysis will be true for any general $f(R)$. The field equations corresponding to the Eqs. (3), (4) in $f(R)$ gravity turns out to be:
$\frac{\dot{a}^{2}}{a^{2}}+\frac{k}{a^{2}}$ $=$ $\frac{8\pi G}{3f^{\prime}(R)}(\rho+\rho_{\rm{eff}})\,$, (24)
$2\frac{\ddot{a}}{a}+\frac{\dot{a}^{2}}{a^{2}}+\frac{k}{a^{2}}$ $=$ $-\frac{8\pi G}{f^{\prime}(R)}(P+P_{\rm{eff}})\,$, (25)
where a prime denotes a derivative with respect to the Ricci scalar, $R$, and
$8\pi G\rho_{\rm{eff}}$ $=$ $\frac{Rf^{\prime}(R)-f(R)}{2}-3H\dot{R}f^{\prime\prime}(R)\equiv\tilde{\rho}_{\rm{eff}}\,$, (26)
$8\pi GP_{\rm{eff}}$ $=$ $\dot{R}^{2}f^{\prime\prime\prime}(R)+2H\dot{R}f^{\prime\prime}(R)+\ddot{R}f^{\prime\prime}(R)-\frac{Rf^{\prime}(R)-f(R)}{2}\equiv\tilde{P}_{\rm{eff}}\,$, (27)
where we have used the Hubble parameter, $H\equiv\dot{a}/a$, to compactly write the above two expressions. The Ricci scalar $R=6[(\ddot{a}/{a})+(\dot{a}^{2}/{a^{2}})+({k}/{a^{2}})]=6[\dot{H}+2H^{2}+(k/a^{2})]$ for the FLRW metric given in Eq. (1). Here $\rho_{\rm{eff}},P_{\rm{eff}}$ are effectively acting like the energy density and the pressure induced by the Ricci curvature of the spacetime. In the above equations $\rho,P$ are the normal energy density and pressure of a barotropic matter fluid present in the FLRW spacetime. We name Eq. (24) as the modified Friedmann equation and Eq. (25) as the modified pressure equation. One can also combine the field equations to write down the modified acceleration equation in the present case. Although we name the relevant field equations as we did before, one must remember that these equations are very different from the corresponding equations in GR as here the modified Friedmann equation is a third order differential equation (apparent from the expression for $\rho_{\rm eff}$), the modified pressure equation and the modified acceleration equation are fourth order differential equations (apparent from the expression for $P_{\rm eff}$).
Remarkably, in the present case if one differentiates the modified Friedmann equation and combines the result with the pressure equation, one gets back the continuity equation for the barotropic fluid, as given in Eq. (6). Consequently, in the present case we have four differential equations of three different orders: one is a third order equation, two are fourth order equations and the other one is a first order equation. The disparity in the degrees of the dynamical equations has increased in the present case. Interestingly, with all these drastic differences with GR, we will show that most of the techniques used to understand the redundancy of the equations are still applicable in the present case and the special role of the modified Friedmann equation remains intact in $f(R)$ gravity.
The above equations show that $f^{\prime}(R)$ modifies the gravitational coupling terms from $G\to G/f^{\prime}(R)$, consequently for gravitational stability of the theory we demand $f^{\prime}(R)>0$. Moreover, it was noted that if $f^{\prime\prime}(R)$ is not positive, one can have another kind of instability of the theory, called the Dolgov-Kawasaki instability [16]. As a consequence of these facts, we will always assume
$f^{\prime}(R)>0\,,\,\,\,\,\,\,\,\,f^{\prime\prime}(R)>0\,$, (28)
in $f(R)$ cosmology. In the field equations of $f(R)$ gravity we have set the cosmological constant, $\Lambda=0$, as $f(R)$ theory can dynamically generate dark energy and we may not explicitly require a cosmological constant in this framework of gravity.
In the present case one can differentiate the modified Friedmann equation with respect to time and combine the result with the continuity equation to obtain the pressure equation. We do not present this derivation as it is a straightforward calculation. We show how one obtains the modified Friedmann constraint from the modified pressure equation and the continuity equation. This calculation will make it clear how the previous concepts can be generalized in the present case. If we choose Eq. (25) and the continuity equation as our basic dynamical equations then we can start to manipulate them in a way done previously. Using the continuity equation we write Eq. (25) as:
$2\frac{\ddot{a}}{a}+\frac{\dot{a}^{2}}{a^{2}}+\frac{k}{a^{2}}+\frac{8\pi G}{f^{\prime}}P_{\rm eff}=\frac{8\pi G}{f^{\prime}}\left(\frac{1}{3}\dot{\rho}\frac{a}{\dot{a}}+\rho\right)\,.$ (29)
Multiplying all the terms of the above equation by $\dot{a}/a$ we get
$2\frac{\ddot{a}\dot{a}}{a^{2}}+\frac{\dot{a}^{3}}{a^{3}}+\frac{k\dot{a}}{a^{3}}+\frac{8\pi G}{f^{\prime}}HP_{\rm eff}=\frac{8\pi G}{3f^{\prime}}\dot{\rho}+\frac{8\pi G}{f^{\prime}}\rho\frac{\dot{a}}{a}\,.$ (30)
The above equation can also be written as
$\frac{d}{dt}\left(\frac{\dot{a}^{2}}{a^{2}}\right)+\frac{3\dot{a}^{3}}{a^{3}}+\frac{k\dot{a}}{a^{3}}+\frac{8\pi G}{f^{\prime}}HP_{\rm eff}=\frac{8\pi G}{3f^{\prime}}\dot{\rho}+\frac{8\pi G}{f^{\prime}}\rho\frac{\dot{a}}{a}\,.$
Subtracting
$3\left(\frac{\dot{a}^{3}}{a^{3}}+\frac{k\dot{a}}{a^{3}}\right)$
from both sides of the last equation, we get
$\frac{d}{dt}\left(\frac{\dot{a}^{2}}{a^{2}}+\frac{k}{a^{2}}\right)+\frac{8\pi G}{f^{\prime}}HP_{\rm eff}=\frac{8\pi G}{3f^{\prime}}\frac{d\rho}{dt}+3\frac{\dot{a}}{a}\left[\frac{8\pi G}{3f^{\prime}}\rho-\left(\frac{\dot{a}^{2}}{a^{2}}+\frac{k}{a^{2}}\right)\right]\,$, (31)
which can also be written as:
$\frac{d}{dt}\left(\frac{\dot{a}^{2}}{a^{2}}+\frac{k}{a^{2}}\right)+\frac{8\pi G}{f^{\prime}}HP_{\rm eff}$ $=$ $\frac{d}{dt}\left[\frac{8\pi G}{3f^{\prime}}(\rho+\rho_{\rm eff})\right]-\frac{8\pi G}{3f^{\prime}}\frac{d\rho_{\rm eff}}{dt}+\frac{8\pi G}{3f^{\prime\,2}}(\rho+\rho_{\rm eff})f^{\prime\prime}\dot{R}$ (32)
- $\frac{8\pi G}{f^{\prime}}H\rho_{\rm eff}+3\frac{\dot{a}}{a}\left[\frac{8\pi G}{3f^{\prime}}(\rho+\rho_{\rm eff})-\left(\frac{\dot{a}^{2}}{a^{2}}+\frac{k}{a^{2}}\right)\right]\,.$
Defining
$Q(t)\equiv\left(\frac{\dot{a}^{2}}{a^{2}}+\frac{k}{a^{2}}\right)-\frac{8\pi G}{3f^{\prime}}(\rho+\rho_{\rm eff})\,$, (33)
we can write Eq. (32) as:
$\frac{dQ}{dt}+3HQ=-\frac{1}{3f^{\prime}}\left[\frac{d\tilde{\rho}_{\rm eff}}{dt}+3H(\tilde{\rho}_{\rm eff}+\tilde{P}_{\rm eff})-3\left\{\frac{8\pi G}{3f^{\prime}}(\rho+\rho_{\rm eff})f^{\prime\prime}\dot{R}\right\}\right]\,.$ (34)
In the present case one can show that:
$\frac{d\tilde{\rho}_{\rm eff}}{dt}+3H(\tilde{\rho}_{\rm eff}+\tilde{P}_{\rm eff})=\dot{R}f^{\prime\prime}\left(\frac{1}{2}R-3\dot{H}-3H^{2}\right)=3\dot{R}f^{\prime\prime}\left(H^{2}+\frac{k}{a^{2}}\right)\,$, (35)
which show that the energy density for the curvature term does not follow a conservation equation. Moreover the above equation reveals that only when $R$ is a constant one may expect that the effective curvature energy density and pressure will satisfy the hydrodynamic equation of continuity. $R$ can turn out to be a constant when $k=0$ and $H$ is a constant, specifying a de Sitter universe. Using the above result we can write Eq. (34) as:
$\frac{dQ}{dt}+3\left(H+\frac{f^{\prime\prime}\dot{R}}{3f^{\prime}}\right)Q=0\,$, (36)
which shows that if $Q=0$ initially, then $Q$ will always be zero. This shows that if we choose the initial conditions to be satisfying the new Friedmann equation then, while working with the new pressure equation and the continuity equation, the new Friedmann equation will always be satisfied.
The above calculation shows that in the present theory, which is very different from GR, we can ask the same questions we asked before. The questions are related to the redundancy of dynamical equations. The redundancy problem ultimately leads to the special status of the modified Friedmann equation. This workout shows that in $f(R)$ cosmology all the initial conditions must have to satisfy the modified Friedmann equation which acts like a constraint equation.
Fortunately, in the present case there exists a particular conformal transformation, which converts the whole problem into a form where the gravitational dynamics is generated by the Einstein equations in GR. Once we convert the problem into a problem of GR, we can then apply our understanding of the initial value problem in GR to unravel the initial value problem in $f(R)$ theories. As because conformal transformations produces GR like dynamics where Einstein equations dictate gravitational dynamics, our old discussions hold in the conformally transformed frame, called the Einstein frame. In this language the system (or the frame) where the dynamics is governed by the original $f(R)$ equations (as above, in particular the Eqs. (24), (25)) is called the Jordan frame. It is seen that our general discussion about the redundancy of cosmological equations and the special role of the Friedmann equation holds true both in the Jordan frame and in the Einstein frame.
7 Conclusion
In this paper we have analyzed the dynamical structure of cosmological evolution in GR. In particular we addressed the issues related to the redundancy of cosmological evolution equations and the relationship of this redundancy with the initial value problem in cosmology. It was then shown that in a modern modified theory of gravity, called $f(R)$ gravity, a similar issue of redundancy of cosmological equations persist. Interestingly, our solution to the redundancy problem in GR sheds important light in solving the redundancy problem in $f(R)$ cosmology. There is a branch of cosmology, called Newtonian cosmology, where one obtains the basic cosmological equations in GR in a dust dominated universe primarily using some adhoc assumptions on the nature of spacetime and semi-Newtonian conceptions [17, 18, 19]. As the basic equations of cosmology in these models remain the same as in GR, one faces the same redundancy problem in these cases and our solution of the redundancy problem works in Newtonian cosmology. It is seen that by focusing our attention on the solvability of the dynamical system we can learn important properties of the Einstein equation in cosmology and unravel the special role of the Bianchi identity. Throughout we have applied a bottom-up approach and it seems that this approach yields interesting results. Our approach will remain relevant, primarily, for two different points. The first point is related to the simplicity and accessibility of the analysis. The present article shows that we can attain some interesting insights about the structure of the cosmological equations starting from very rudimentary assumptions. The second point is related to the fact that the results about the redundancy problem in cosmology in GR can be generalized to more modern variants of theory of gravity, as $f(R)$ gravity.